# Personalisation via Dynamic Policy Fusion

### v2

@@ Abstract
Deep reinforcement learning (RL) policies, although optimal in terms of task rewards, may not align with the personal preferences of human users. To ensure this alignment, a naïve solution would be to retrain the agent using a reward function that encodes the user’s specific preferences. However, such a reward function is typically not readily available, and as such, retraining the agent from scratch can be prohibitively expensive. We propose a more practical approach - to adapt the already trained policy to user-specific needs with the help of human feedback. To this end, we infer the user’s intent through trajectory-level feedback and combine it with the trained task policy via a theoretically grounded dynamic policy fusion approach. As our approach collects human feedback on the very same trajectories used to learn the task policy, it does not require any additional interactions with the environment, making it a zero-shot approach. We empirically demonstrate in a number of environments that our proposed dynamic policy fusion approach consistently achieves the intended task while simultaneously adhering to user-specific needs.
@@
Reinforcement Learning, Personalisation, Dynamic Policy Fusion
@@ 1 Introduction
Reinforcement learning (RL) has demonstrated its effectiveness in various real-world scenarios, including computer games [16], inventory management, autonomous driving, robotics [8], healthcare [30], recommendation systems [25] etc. In RL, the learning agent typically learns to maximise its reward by learning an optimal task policy through interactions with the environment. However, in certain scenarios, a user may desire a policy that subtly deviates from this optimal policy to accommodate their personal preference or style. For example, in a navigation setting, users may want to avoid toll roads, or they may prefer to drive along a scenic coastal route, even if this translates to slightly longer travel times compared to the route determined to be the optimal one by an RL agent trained to minimise travel time.
@@
We note that the problem of respecting additional user-specific objectives could also be formulated as a multi-objective reinforcement learning [22, 9, 15] problem, although such approaches are applicable only when rewards corresponding to each objective are available during learning. As such, the present work aims to deal with how one can adapt a trained policy after it has learned an initial task. Overall, the contributions of this work can be summarised as follows:
  • A technique to achieve zero-shot personalisation by adapting an already trained task policy without additional environmental interactions.
  • Theoretical analysis establishing the boundedness of the personalised policy with respect to the task policy.
  • Demonstration of the failure of static policy fusion methods, motivating the need for dynamic fusion techniques.
  • A modulation mechanism for dynamic policy fusion to balance task performance with user preference.
@@ 3.2 RUDDER
We use the RUDDER framework to learn the Q-values corresponding to human scores and later convert the Q-values to policies as described in Sections 4.1 and 4.2. The densified reward as in (1) is used to modulate the influence of the policy as discussed in Section 4.5. For more details of the training of LSTM, please refer to Section 0.B of the supplementary.
@@ 4 Methodology
We construct our proposed zero-shot approach for personalisation by first inferring the intent-specific policy via trajectory-level human feedback using RUDDER, followed by a dynamic policy fusion mechanism that automatically maintains a balance between the inferred policy and a trained task-specific policy.
@@
Figure 1: Summary of our method. Trajectory-level training data used to train the task policy is labelled by a human user and an LSTM model is employed to identify the intent of the human. The output from the LSTM is converted to an intent-specific policy which is then dynamically fused by modulating the temperature parameter of the intent-specific policy (Discussed in Section 4.5) with the task-specific policy to obtain the personalised policy. The blue shaded block is the task policy training and the green dotted block is the dynamic policy fusion.
@@ 4.1 Learning Human Intent using LSTM
To learn the intent-specific policy, as previously described, we leverage the LSTM-based approach of RUDDER. We train this LSTM using human trajectory-level feedback on the same training data used to learn the task-specific policy. Hence, no additional interaction data is required. To get the human feedback, we simulated the human. The feedback score is the number of times the human preference is met within a trajectory (Refer to supplementary Section 0.A for more details).
At each time step, the state-action vector is fed into the LSTM units. In the case of image inputs, we pre-train a Variational Autoencoder (VAE) to reduce the dimensionality of the state. We use FiLM [20] to modulate the state vector with the action vector. In [20], a FiLM network conditions the feature map of the neural network depending on another input signal. Here, we condition the state feature with the action vector and train the LSTM using the human feedback data as training labels. With this training setup, (More details in supplementary Section 0.B and 0.C) the model approximates the Q-value, which is then converted to the intent-specific policy as described in the next section.
@@ 4.2 Policy Construction
where $T_{\phi}$ is the temperature corresponding to task-specific policy and $T_{\psi}$ is the temperature corresponding to intent-specific policy. With policies constructed from Q-values, these constituent policies are fused together to produce the personalised policy
@@ 4.3 Policy Fusion
Definition 1 (Invariant policies)
@@
The above lemma captures the invariability of the proposed fused method when $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$. However, if the deviation between the Q-values of the intent-specific policy and the task-specific policy is bounded (given the deviation of the temperature of both policies is bounded), then the divergence between the task-specific policy and the personalised policy can be bounded as stated in Theorem 4.1. This theorem constraints the proposed fusion method to arbitrarily deviate from the task policy.
Theorem 4.1
@@
where $Z$ is the normalising factor of $\pi_{f}$, $Q^{*}=argmax_{a\in\mathcal{A}}Q(s,a)$ and $\zeta=\dfrac{h(Q^{\prime},T_{\psi})}{h(Q,T_{\phi})}$, here $h(Q,T)=\sum_{a}exp\left(\dfrac{Q}{T}\right)$.
Proof
@@
$=\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}+QT_{\phi}-QT_{\phi}}{T_{\phi}T_{\psi}}\right)+\dfrac{1}{2}\log\left(\zeta\right)$
@@
The above theorem establishes the boundedness of the divergence between the task-specific policy and the personalised policy. Nevertheless, we can consider other choices for policy fusion, such as the weighted average or the product of the two policies. However, the former does not operate on the common support, and the latter does not satisfy the invariability constraint, even when the policies are exactly the same, as evidenced in the Lemma 2. For a detailed analysis, we refer the reader to the supplementary Section 0.E.
Lemma 2
@@
Proof
$KL\left(\pi_{\phi}|\pi_{\phi}\pi_{\psi}\right)=$ $\sum_{a}\pi_{\phi}\log\left(\dfrac{\pi_{\phi}Z}{\pi_{\phi}\pi_{\psi}}\right)$
@@
This expression is valid only if $\pi_{\phi}$ is a random policy, which is a contradiction.
@@ 4.4 Pitfalls of Static Fusion
4.5 Dynamically Modulating Policy Dominance
To modulate $T_{\psi}$, we first compute the shifted rewards as follows:
$r^{\prime}=r-mean(r),$ (5)
@@ 5 Experiments
We demonstrate the effectiveness of our approach in our custom 2D Navigation environment, Highway [highwayenv], Pong [26] depicted in Figure 3. Specifically, we conduct personalisation experiments in three scenarios: Avoidance, Preference and Mixed  [^1] Code for our implementation can be found here. For Pong, we only consider the Preference scenario. To learn the intent-specific policy, we sampled 2000 trajectories from saved trajectory data. Now we describe the environments as follows:
@@
Env Mode Human objective Task objective 2D Navigation Preference Prefer a region Move to the target Avoidance Avoid a region Mixed Prefer a region and avoid another Pong Preference Prefer certain paddle positions Win the game Highway Preference Prefer a lane Move as fast as possible Avoidance Avoid a lane without colliding Mixed Prefer a lane and avoid another with other vehicles
@@
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$ Preference DQN $0.085\pm 0.005$ $1.000\pm 0.000$ RUDDER $6.546\pm 0.342$ $0.006\pm 0.005$ MORL $9.575\pm 0.021$ $000.000\pm 0$ 0.000 Static $3.788\pm 0.452$ $0.602\pm 0.048$ Dynamic (Ours) $1.459\pm 0.140$ $1.000\pm 0.000$ Mixed DQN $0.050\pm 0.003$ $0.183\pm 0.006$ $1.000\pm 0.000$ RUDDER $2.790\pm 0.311$ $000.000\pm 0$ 0.000 $0.007\pm 0.007$ MORL $0.192\pm 0.022$ $0.119\pm 0.016$ $0.396\pm 0.035$ Static $1.116\pm 0.318$ $000.000\pm 0$ 0.000 $0.894\pm 0.035$ Dynamic (Ours) $0.268\pm 0.058$ $000.000\pm 0$ 0.000 $1.000\pm 0.000$ Avoidance DQN - $0.094\pm 0.006$ $1.000\pm 0.000$ RUDDER - $000.000\pm 0$ 0.000 $0.292\pm 0.055$ MORL - $0.026\pm 0.021$ $0.894\pm 0.099$ Static - $0.005\pm 0.005$ $1.000\pm 0.000$ Dynamic (Ours) - $0.006\pm 0.005$ $1.000\pm 0.000$
@@ 5.1 Results
We further examine the properties of dynamic fusion in the Highway and Pong environments in Tables 3 and 4 respectively with the results averaged over 10 seeds each with 20 episodes. Here, DQN and RUDDER perform as expected, prioritising the task and human objectives respectively. In Highway, as seen in Table 3, DQN maximises the score and minimises hits (collisions), whereas RUDDER exhibits behaviours avoiding the undesired (first) lane and favouring the desired (third) lane. Similarly, as indicated in Table 4 in Pong, while DQN maximises the task score at the cost of seldom visiting the desired region, RUDDER causes the agent to remain in the preferred region, consistent with human preference (but incurring low scores). When compared to the dynamic policy fusion, the performance of the MORL baseline falls short of achieving both objectives. This is evident from the higher frequency of undesired lane visits in both the Avoidance and Mixed modes, and the lower frequency of desired lane visits in the Preference and Mixed modes. These results are detailed in Table 3. A similar trend is observed in Pong, where the MORL baseline not only failed to score but also spent less time in the preferred region compared to our method. These results indicate that dynamic policy fusion, while sometimes incurring a small reduction in the optimal score (achieved by DQN), enables the agent to adhere to the required preference (be it in the Avoiding, Preference or Mixed modes) while completing the primary task at hand. Additional experiments on different $\alpha$ in the Highway environment are provided in supplementary Section 0.G.
Mode Method Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$ Avoidance DQN - $10.59\pm 0.96$ $0.11\pm 0.01$ $39.59\pm 0.46$ RUDDER - $00.00\pm 0$ 0.00 $0.73\pm 0.11$ $18.43\pm 2.54$ MORL - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$ Dynamic (Ours) - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$ Preference DQN $10.9\pm 0.49$ $0.07\pm 0.02$ $40.25\pm 0.64$ RUDDER $31.84\pm 3.45$ $0.30\pm 0.06$ $28.45\pm 1.65$ MORL $11.58\pm 1.12$ $0.07\pm 0.01$ $38.14\pm 0.30$ Dynamic (Ours) $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$ Mixed DQN $11.83\pm 0.64$ $9.59\pm 0.82$ $0.08\pm 0.02$ $40.06\pm 0.79$ RUDDER $26.26\pm 5.90$ $00.00\pm 0$ 0.00 $0.50\pm 0.12$ $24.67\pm 3.00$ MORL $12.61\pm 1.20$ $12.36\pm 1.58$ $0.07\pm 0.01$ $36.97\pm 0.59$ Dynamic (Ours) $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
@@
Method % Desired region $\uparrow$ Score $\uparrow$ DQN $8.00\pm 0.63$ $0.57\pm 0.01$ RUDDER $89.99\pm 0.14$ $00.00\pm 0$ 0.00 MORL $41.68\pm 0.02$ $00.00\pm 0$ 0.00 Dynamic (Ours) $55.60\pm 4.48$ $0.48\pm 0.02$
@@
Mode $\eta$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$ Avoidance 0 - $0.28\pm 0.08$ $0.06\pm 0.01$ $39.83\pm 0.59$ 1 - $0.35\pm 0.12$ $0.08\pm 0.02$ $38.96\pm 0.87$ 2 - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$ Preference 0 $25.14\pm 1.44$ $0.05\pm 0.02$ $40.02\pm 0.61$ 1 $24.90\pm 1.52$ $0.10\pm 0.03$ $38.84\pm 0.84$ 2 $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$ Mixed 0 $21.52\pm 2.00$ $0.85\pm 0.33$ $0.12\pm 0.01$ $38.50\pm 0.43$ 1 $23.96\pm 1.67$ $0.56\pm 0.13$ $0.07\pm 0.01$ $39.56\pm 0.45$ 2 $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
@@
Effect of $\eta$: In Table 5, we examine the effect of $\eta$ on the performance of an agent under dynamic fusion. In theory, a higher $\eta$ would make it more likely for $T_{\psi}$ to remain low (refer to Figure 2), and thus the intent-specific policy becomes sharper. So while fusing the policies together, the obtained personalised policy become dominated by intent-specific policy. Our observations in Table 3 are in agreement with this behaviour, as in all preference modes, as $\eta$ increases, the desired lane visit increases and the undesired lane visit decreases, suggesting that personalised policy is dominated by the intent-specific policy component. In all of our experiments, we choose $\eta=0$. However, an ideal choice of $\eta$ may deviate from this value in other environments.
Effect of $T_{max}$: The parameter $T_{max}$ represents the maximum temperature in Equation (6). Theoretically, as $T_{max}$ is increased the intent-specific policy should weaken. The observation in Table 6 supports this notion that when $T_{max}$ is increased, the desired lane visit decreases and the undesired lane visit increases suggesting a declining influence of intent-specific policy.

### v3

@@ Abstract
Deep reinforcement learning (RL) policies, although optimal in terms of task rewards, may not align with the personal preferences of human users. To ensure this alignment, a naïve solution would be to retrain the agent using a reward function that encodes the user’s specific preferences. However, such a reward function is typically not readily available, and as such, retraining the agent from scratch can be prohibitively expensive. We propose a more practical approach - modify the already trained policy to accommodate the user-specific needs with the help of human feedback. To this end, we infer the user’s intent through trajectory-level feedback and combine it with the trained task policy via a theoretically grounded dynamic policy fusion approach. As our approach collects human feedback on the very same trajectories used to learn the task policy, it does not require any additional interactions with the environment, making it a zero-shot approach in terms of environment interactions. We empirically demonstrate in a number of environments that our proposed dynamic policy fusion approach consistently achieves the intended task while simultaneously adhering to user-specific needs.
@@
Reinforcement Learning, User Alignment, Dynamic Policy Fusion, Human in the Loop.
@@ 1 Introduction
Reinforcement learning (RL) has demonstrated its effectiveness in various real-world scenarios, including computer games [16], inventory management, autonomous driving, robotics [8], healthcare [31], recommendation systems [26] etc. In RL, the learning agent typically learns to maximise its reward by learning an optimal task policy through interactions with the environment. However, in certain scenarios, a user may desire a policy that subtly deviates from this optimal policy to accommodate their personal preference or style. For example, in a navigation setting, users may prefer to take quieter, less congested routes, even if this means slightly longer travel times compared to the route determined to be the optimal one by a reinforcement learning (RL) agent trained to minimize travel time. This preference allows users to enjoy a more relaxed and stress-free journey.
@@
We note that the problem of respecting additional user-specific objectives could also be formulated as a multi-objective reinforcement learning [22, 9, 15] problem, although such approaches are applicable only when rewards corresponding to each objective are available during learning. As such, the present work aims to deal with how one can adopt a trained policy after one has learned an initial task.
Some of our early work in this direction has been reported in [25] and our current work provides a comprehensive presentation including improvements in the method, new theoretical analysis to firmly ground the main intuition behind the policy fusion method, and extensive evaluation across diverse environments.
@@ 3.2 RUDDER
We use the RUDDER framework to learn the Q-values corresponding to human scores and later convert the Q-values to policies as described in Sections 4.1 and 4.2. The densified reward as in (1) is used to modulate the influence of the policy as discussed in Section 4.5. For more details of the training of LSTM, please refer to Section B of the supplementary.
@@ 4 Methodology
We construct our proposed approach for personalisation by first inferring the intent-specific policy via trajectory-level human feedback using RUDDER, followed by a dynamic policy fusion mechanism that automatically maintains a balance between the inferred policy and a trained task-specific policy.
@@
Figure 1: Summary of our method. Trajectory-level training data used to train the task policy is labelled by a human user and an LSTM model is employed to identify the intent of the human. The output from the LSTM is converted to an intent-specific policy which is then dynamically fused by modulating the temperature parameter of the intent-specific policy (Discussed in Section 4.5) with the task-specific policy to obtain the personalised policy. The blue shaded block is the task policy training and the green dotted block is the dynamic policy fusion. (Image taken from previous study [25])
@@ 4.1 Learning Human Intent using LSTM
To learn the intent-specific policy, as previously described, we leverage the LSTM-based approach of RUDDER. We train this LSTM using human trajectory-level feedback on the same training data used to learn the task-specific policy. Hence, no additional interaction data is required. To get the human feedback, we simulated the human. The feedback score is the number of times the human preference is met within a trajectory (Refer to supplementary Section A for more details).
At each time step, the state-action vector is fed into the LSTM units. In the case of image inputs, we pre-train a Variational Autoencoder (VAE) to reduce the dimensionality of the state. We use FiLM [20] to modulate the state vector with the action vector. In [20], a FiLM network conditions the feature map of the neural network depending on another input signal. Here, we condition the state feature with the action vector and train the LSTM using the human feedback data as training labels. With this training setup, (More details in supplementary Section B and C) the model approximates the Q-value, which is then converted to the intent-specific policy as described in the next section.
@@ 4.2 Policy Construction
where $T_{\phi}$ is the temperature corresponding to task-specific policy and $T_{\psi}$ is the temperature corresponding to intent-specific policy. With policies constructed from Q-values, these constituent policies are fused together to produce the personalised policy.
@@ 4.3 Policy Fusion
Definition 1 (Invariant policies).
@@
The above lemma captures the invariability of the proposed fused method when $\pi_{\phi}(a|s)=\pi_{\psi}(a|s)$. However, if the deviation between the Q-values of the intent-specific policy and the task-specific policy is bounded (given the deviation of the temperature of both policies is bounded), then the divergence between the task-specific policy and the personalised policy can be bounded as stated in Theorem 1. This theorem constraints the proposed fusion method to arbitrarily deviate from the task policy.
Theorem 1.
@@
where $Z$ is the normalising factor of $\pi_{f}$, $Q^{*}=max_{a\in\mathcal{A}}Q(s,a)$ and $\zeta=\dfrac{h(Q^{\prime},T_{\psi})}{h(Q,T_{\phi})}$, here $h(Q,T)=\sum_{a}exp\left(\dfrac{Q}{T}\right)$.
Proof.
@@
$=\log\left(Z\right)+\sum_{a}\dfrac{\pi_{\phi}}{2}\left(\dfrac{QT_{\psi}-Q^{\prime}T_{\phi}+QT_{\phi}-QT_{\phi}}{T_{\phi}T_{\psi}}\right)$
$+\dfrac{1}{2}\log\left(\zeta\right)$
@@
The above theorem establishes the boundedness of the divergence between the task-specific policy and the personalised policy. Nevertheless, we can consider other choices for policy fusion as studied in the previous work [24], such as the weighted average or the product of the two policies. However, the former does not operate on the common support, and the latter does not satisfy the invariability constraint, even when the policies are exactly the same, as evidenced in the Lemma 2. For a detailed analysis, we refer the reader to the supplementary Section E.
Lemma 2.
@@
Proof.
$KL\left(\pi_{\phi}|\tfrac{\pi_{\phi}\pi_{\psi}}{Z}\right)=$ $\sum_{a}\pi_{\phi}\log\left(\dfrac{\pi_{\phi}Z}{\pi_{\phi}\pi_{\psi}}\right)$
@@
The Jensen’s Inequality states that $E_{\pi_{\phi}}\left(\log\left(\pi_{\psi}\right)\right)\leq\log\left(E_{\pi_{\phi}}\left(\pi_{\psi}\right)\right)$ and this expression is an equality only if $\pi_{\psi}$ is a random policy, which is a contradiction. ∎
@@ 4.4 Pitfalls of Static Fusion
4.5 Dynamically Modulating Policy Dominance
To modulate $T_{\psi}$, we first compute the shifted rewards as follows:
$r^{\prime}=r-mean(r),$ (5)
Env Mode Human objective Task objective
2D Navigation Preference Prefer a region Move to the target
Avoidance Avoid a region
Mixed Prefer a region and avoid another
Pong Preference Prefer certain paddle positions Win the game
Highway Preference Prefer a lane Move as fast as possible
Avoidance Avoid a lane without colliding
Mixed Prefer a lane and avoid another with other vehicles
@@ 5 Experiments
We demonstrate the effectiveness of our approach in our custom 2D Navigation environment, Highway [13], Pong [27] depicted in Figure 3. Specifically, we conduct personalisation experiments in three scenarios: Avoidance, Preference and Mixed  [^1] Code for our implementation can be found here. For Pong, we only consider the Preference scenario. To learn the intent-specific policy, we sampled 2000 trajectories from saved trajectory data. Now we describe the environments as follows:
@@
Mode Method Desired region $\uparrow$ Undesired region $\downarrow$ Score $\uparrow$
Preference DQN $0.085\pm 0.005$ $1.000\pm 0.000$
RUDDER $6.546\pm 0.342$ $0.006\pm 0.005$
MORL $9.575\pm 0.021$ $000.000\pm 0$ 0.000
Static $3.788\pm 0.452$ $0.602\pm 0.048$
Dynamic (Ours) $1.459\pm 0.140$ $1.000\pm 0.000$
Mixed DQN $0.050\pm 0.003$ $0.183\pm 0.006$ $1.000\pm 0.000$
RUDDER $2.790\pm 0.311$ $000.000\pm 0$ 0.000 $0.007\pm 0.007$
MORL $0.192\pm 0.022$ $0.119\pm 0.016$ $0.396\pm 0.035$
Static $1.116\pm 0.318$ $000.000\pm 0$ 0.000 $0.894\pm 0.035$
Dynamic (Ours) $0.268\pm 0.058$ $000.000\pm 0$ 0.000 $1.000\pm 0.000$
Avoidance DQN - $0.094\pm 0.006$ $1.000\pm 0.000$
RUDDER - $000.000\pm 0$ 0.000 $0.292\pm 0.055$
MORL - $0.026\pm 0.021$ $0.894\pm 0.099$
Static - $0.005\pm 0.005$ $1.000\pm 0.000$
Dynamic (Ours) - $0.006\pm 0.005$ $1.000\pm 0.000$
Mode Method Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$
Avoidance DQN - $10.59\pm 0.96$ $0.11\pm 0.01$ $39.59\pm 0.46$
RUDDER - $00.00\pm 0$ 0.00 $0.73\pm 0.11$ $18.43\pm 2.54$
MORL - $5.80\pm 1.07$ $0.14\pm 0.23$ $37.44\pm 0.52$
Dynamic (Ours) - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$
Preference DQN $10.9\pm 0.49$ $0.07\pm 0.02$ $40.25\pm 0.64$
RUDDER $31.84\pm 3.45$ $0.30\pm 0.06$ $28.45\pm 1.65$
MORL $11.58\pm 1.12$ $0.07\pm 0.01$ $38.14\pm 0.30$
Dynamic (Ours) $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$
Mixed DQN $11.83\pm 0.64$ $9.59\pm 0.82$ $0.08\pm 0.02$ $40.06\pm 0.79$
RUDDER $26.26\pm 5.90$ $00.00\pm 0$ 0.00 $0.50\pm 0.12$ $24.67\pm 3.00$
MORL $12.61\pm 1.20$ $12.36\pm 1.58$ $0.07\pm 0.01$ $36.97\pm 0.59$
Dynamic (Ours) $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
@@ 5.1 Results
We further examine the properties of dynamic fusion in the Highway and Pong environments in Tables 3 and 4 respectively with the results averaged over 10 seeds each with 20 episodes. Here, DQN and RUDDER perform as expected, prioritising the task and human objectives respectively. In Highway, as seen in Table 3, DQN maximises the score and minimises hits (collisions), whereas RUDDER exhibits behaviours avoiding the undesired (first) lane and favouring the desired (third) lane. Similarly, as indicated in Table 4 in Pong, while DQN maximises the task score at the cost of seldom visiting the desired region, RUDDER causes the agent to remain in the preferred region, consistent with human preference (but incurring low scores). When compared to the dynamic policy fusion, the performance of the MORL baseline falls short of achieving both objectives. This is evident from the higher frequency of undesired lane visits in both the Avoidance and Mixed modes, and the lower frequency of desired lane visits in the Preference and Mixed modes. These results are detailed in Table 3. A similar trend is observed in Pong, where the MORL baseline not only failed to score but also spent less time in the preferred region compared to our method. These results indicate that dynamic policy fusion, while sometimes incurring a small reduction in the optimal score (achieved by DQN), enables the agent to adhere to the required preference (be it in the Avoiding, Preference or Mixed modes) while completing the primary task at hand. Additional experiments on different $\alpha$ in the Highway environment are provided in supplementary Section G.
@@
Method % Desired region $\uparrow$ Score $\uparrow$
DQN $8.00\pm 0.63$ $0.57\pm 0.01$
RUDDER $89.99\pm 0.14$ $00.00\pm 0$ 0.00
MORL $41.68\pm 0.02$ $00.00\pm 0$ 0.00
Dynamic (Ours) $55.60\pm 4.48$ $0.48\pm 0.02$
@@
Mode $\eta$ Desired lane $\uparrow$ Undesired lane $\downarrow$ Hits $\downarrow$ Score $\uparrow$
Avoidance 0 - $0.28\pm 0.08$ $0.06\pm 0.01$ $39.83\pm 0.59$
1 - $0.35\pm 0.12$ $0.08\pm 0.02$ $38.96\pm 0.87$
2 - $0.19\pm 0.11$ $0.09\pm 0.02$ $38.80\pm 0.51$
Preference 0 $25.14\pm 1.44$ $0.05\pm 0.02$ $40.02\pm 0.61$
1 $24.90\pm 1.52$ $0.10\pm 0.03$ $38.84\pm 0.84$
2 $29.91\pm 1.38$ $0.06\pm 0.01$ $39.27\pm 0.59$
Mixed 0 $21.52\pm 2.00$ $0.85\pm 0.33$ $0.12\pm 0.01$ $38.50\pm 0.43$
1 $23.96\pm 1.67$ $0.56\pm 0.13$ $0.07\pm 0.01$ $39.56\pm 0.45$
2 $25.21\pm 2.55$ $0.27\pm 0.06$ $0.07\pm 0.04$ $39.15\pm 0.49$
@@
10 $19.21\pm 1.42$ $2.02\pm 0.43$ $0.06\pm 0.02$ $40.12\pm 0.59$
15 $17.60\pm 0.90$ $3.42\pm 0.43$ $0.09\pm 0.02$ $39.64\pm 0.45$
25 $9.53\pm 0.82$ $4.65\pm 0.52$ $0.10\pm 0.02$ $39.51\pm 0.53$
35 $14.36\pm 0.65$ $5.32\pm 0.58$ $0.10\pm 0.02$ $40.03\pm 0.59$
Effect of $\eta$: In Table 5, we examine the effect of $\eta$ on the performance of an agent under dynamic fusion. In theory, a higher $\eta$ would make it more likely for $T_{\psi}$ to remain low (refer to Figure 2), and thus the intent-specific policy becomes sharper. So while fusing the policies together, the obtained personalised policy becomes dominated by intent-specific policy. Our observations in Table 3 are in agreement with this behaviour, as in all preference modes, as $\eta$ increases, the desired lane visit increases and the undesired lane visit decreases, suggesting that personalised policy is dominated by the intent-specific policy component. In all of our experiments, we choose $\eta=0$. However, an ideal choice of $\eta$ may deviate from this value in other environments.
Effect of $T_{max}$: The parameter $T_{max}$ represents the maximum temperature in Equation (6). Theoretically, as $T_{max}$ is increased, the intent-specific policy should weaken. The observation in Table 6 supports this notion that when $T_{max}$ is increased, the desired lane visit decreases and the undesired lane visit increases suggesting a declining influence of intent-specific policy.